Classical Algebra (Math 326) Written Assignment

due February 11, 2000

Do one of the following:

A. Submit properly written solutions to the following problems in chapter 7 of the text by Childs:

7 D: 1, 2, & 3.

OR

B. Do the following:

1. Write a computer program that lists all primes up to a given value N that is the program’s “argument”.

2. Make the source code for the program available in your on-campus web site along with a file containing the output from your program in the case when the given number N is your “key”, as defined below, anchoring these two documents from a properly prepared web page that explains briefly what you have done and introduces the two exhibits. The URL for the web page should be of the form

                 http://www.albany.edu/~login-id/math326/assgt1.html

unless you present a reasonable excuse for submission in some other form.

Your key may be either the last 4 non-zero digits in your student-id number OR the number constructed from your date of birth by using the first 4 non-zero digits taken in order from the 8 digits obtained by putting that date in the form DDMMYYYY.

3. Submit a written record of your work with

(a) a printed copy of the source code for your program.

(b) your key.

(c) the last 5 primes before your key.

(d) the exact URL for the web page (as above).

Keep the URL in force through the last day of examinations this semester.

Document network location:

DVI: http://math.albany.edu:8000/math/pers/hammond/course/mat326s2000/assgt/cla000202.dvi
HTML: http://math.albany.edu:8000/math/pers/hammond/course/mat326s2000/assgt/cla000202.html